# A Scene Representation for Online Spatial Sonification

### v1

@@ Abstract
Robotic perception is emerging as a crucial technology for navigation aids, particularly benefiting individuals with visual impairments through sonification. This paper presents a novel mapping framework that accurately represents spatial geometry for sonification, transforming physical spaces into auditory experiences. By leveraging depth sensors, we convert incrementally built 3D scenes into a compact 360-degree representation based on angular and distance information, aligning with human auditory perception. Our proposed mapping framework utilises a sensor-centric structure, maintaining 2D circular or 3D cylindrical representations, and employs the VDB-GPDF for efficient online mapping. We introduce two sonification modes—circular ranging and circular ranging of objects—along with real-time user control over auditory filters. Incorporating binaural room impulse responses, our framework provides perceptually robust auditory feedback. Quantitative and qualitative evaluations demonstrate superior performance in accuracy, coverage, and timing compared to existing approaches, with effective handling of dynamic objects. The accompanying video showcases the practical application of spatial sonification in room-like environments.
Keywords: Spatial sonification, Scene Representation, Gaussian Process Distance Field.
@@ I Introduction
Robotic perception has the potential to see the environment for human beings. Over the last two decades, robotic sensors such as cameras or LiDARs have served as eyes for robots and have become popular for gathering information from unknown environments. For people living with vision disability, auditory perception is required and could be fulfilled by transforming the information accumulated by the robotic perception via sonification. One of the key steps in representing an environment for sonification is to convert spatial data into an organised format that can be understood through sound. This process facilitates the transformation of physical spaces into auditory experiences, allowing users to understand and navigate through sound. The primary goal of this work is to build a robust framework that can precisely represent an environment’s spatial properties such that it can serve as a basis for sonification.
@@
The ability to incrementally build and maintain a compact representation through sensors will enable real-time auditory feedback to support spatial awareness and environment interaction, especially for visually impaired individuals. As we want our representation to cater towards sonification for humans, we take motivation from how human beings perceive an environment through sound. Thus, we transform a 3D scene captured with depth data into a 360-degree singular representation, based on the angular and distance information. This approach naturally aligns with human auditory perception while simplifying information by focusing on the most relevant spatial points. To achieve this, we propose a mapping framework, based on a sensor-centric organised structure that maintains 2D circular and/or 3D cylindrical rasterised representations (see Fig. 1). These structures capture the geometry of the 3D environment and offer flexibility in how spatial points are sonified by leveraging VDB-GPDF [1], an online efficient mapping framework based on Gaussian Process (GP) distance and gradient fields and a fast-access VDB data structure (Volumetric, Dynamic grid that shares several characteristics with B+trees [2, 3]), enabling a sonifiable online incrementally-built scene representation for indoor and outdoor scenarios.
Figure 1: Our sensor-centric representation for spatial sonification. a) 2D circular representation and b) 3D cylindrical representation for the Cow and Lady dataset.
Moreover, we demonstrate two modes of operation, a circular ranging and a circular ranging of objects sonification modes, which sonify the distance to objects for various directions around the circle. We also introduce various filtering operations to extract information from the proposed 2D circular representation. By incorporating binaural room impulse responses (BRIRs), our mapping framework offers perceptually robust auditory feedback and enhances cognitive mapping. This combination allows for customised sonification of both near-field and far-field objects.
@@
We evaluate quantitatively the performance of the proposed mapping framework in terms of accuracy, coverage and timing, demonstrating the ability to convey the required spatial information better than naively using the current sensor information. Moreover, we show qualitative performance in the presence of dynamic objects and benchmarked with the information provided by distance field mapping. Finally, the submitted video presents the spatial sonification for room-like environments.
@@ II Literature Review
Several studies have explored sonification from vision-based representations. While earlier works focused mostly on 2D vision-to-audio transformations [4, 5], more recent approaches utilise 3D vision-based sensors [6, 7]. Although 3D representations generally offer improved accuracy for localisation tasks, 2D systems can achieve comparable performance for navigation with sonification [8]. Some works have directly used 3D raw data for environment representation [9, 10]. In [9], the authors presented a real-time auditory feedback system that assists blind users in obstacle navigation using a depth camera. The system processes depth data to detect obstacles and sonifies this information. However, using depth data directly can be complex for turning 3D information into clear auditory cues. To tackle this, we propose a 360-degree circular/cylindrical rasterisation method that projects the 3D scenes onto a minimalist geometric structure.
@@
More recently, there has been a growing trend towards leveraging machine learning techniques, such as generative models, to enhance environment representation for spatial sonification. These methods allow for more adaptive and context-aware sonification systems, improving the overall user experience. In [11], the authors introduced “SEE-2-SOUND,” a zero-shot framework. They generated spatial audio from visual inputs such as images, GIFs, and videos without pre-training. A similar task was performed in [12], with additional use of a large language model to generate context-aware synchronised audio-visual content based on the event. While these systems offer adaptability, they often require substantial computational resources. This introduces substantial latency in the sonification process. In contrast, our method provides a lightweight, direct mapping of spatial data to auditory cues. This ensures real-time responsiveness while preserving essential spatial details and the ability to be deployed on modest embedded devices.
@@
There are different techniques to sonify a virtual environment, Gao et al. showed quantitative evidence of how different sonification strategies can influence user perception and behaviour [13]. They compared four elevation sonification mapping strategies and found that binary relative mapping performs better than the others, and in combination with azimuth sonification, it gives clear directional information generating more accurate and efficient task completion. Keeping this in mind, our proposed representation technique is designed to effectively capture both angular and distance information. Thus, ensuring optimal translation of spatial data into auditory cues for the sonification process.
@@
There have been many works on sonification that are dedicated towards visually impaired persons[14, 15]. The type of sound used for sonification also varies depending on the task and the type of information it carries. While discrete sounds are good for simpler information and objects that are near the user, continuous sounds are more suited to deliver complex elements in the scene [16]. Schwartz et al. proposed a mobile application that can reconstruct and sonify a scene to guide users with visual impairment in real-time [15]. The work in [17] built a 3D sonification framework to sonify images from video sequences to send the user a predefined sonic sensorial description through bone-conduction headphones. They combined YOLOV3 [18] for object detection with a probabilistic occupancy map from OctoMap [19] to detect environmental anomalies. These works primarily focus on sonifying immediate surroundings or predefined objects with limited adaptability. They often lack dynamic user control over the sonification process and do not offer full 360-degree spatial coverage, focusing mostly on near-field objects. Our mapping framework, in contrast, is flexible, robust and offers greater adaptability as it conveys spatial information for near-field and far-field.
@@ III Preliminaries
III-A VDB-GPDF Mapping
@@
For testing, a set of query points is generated by ray-casting from the sensor origin through the voxels in the current reference frame. Each test point is used to query the L-GPDF for the distance field and variance estimates. The queried values from L-GPDF are then fused into a global VDB grid map by updating the voxel distances using the weighted sum method. Following the fusion of the distances and surface properties, the marching cubes algorithm is used to reconstruct a dense surface by identifying the active voxels in the global VDB. The Global Gaussian Process Distance Field (G-GPDF) is made up of the zero-crossing points from active leaf nodes. At this stage, queries for the G-GPDF are computed by averaging the inferred distances from neighboring GP nodes. For further details refer to [1].
IV Methodology
We present an incrementally built scene representation that aims to facilitate spatial sonification of an unknown environment. The proposed representation is based on VDB-GPDF and expanded to cater sonification requirements for human sensory augmentation. As mentioned above, the 3D surface is extracted by applying the Marching Cubes algorithm over the G-GPDF model. propose to simplify this 3D mesh into two distinct rasterisations. First, we squash the 3D surface mesh points into their respective 2D coordinates, and then, these are used to create a 2D circular structure. We also do a simplified 3D cylindrical rasterisation while retaining the full 3D spatial structure. While the 2D circular rasterisation simplifies the 3D environment by projecting the mesh points onto a single 2D circle, the 3D cylindrical rasterisation represents both the horizontal and vertical points on the surface, capturing not only the azimuthal angles but also the height (elevation) of each point. To show the potential of the representation to sonify a 3D environment we develop various strategies for spatial audio feedback, facilitating intuitive sonification with auditory cues.
@@ IV Methodology
Spatial audio feedback requires accurate but simple information, for this purpose, we propose to represent a dense 3D map of the environment as a 2D circular sensor-centric representation. To ensure our representation aligns with human navigation requirements, we segment ground and non-ground points. Any point that is above the ground and in the height range of a human and can potentially be an obstacle is classified as a non-ground point. Then, we project all these 3D non-ground point cloud or mesh onto a 2D plane. On this plane, every 3D point is reduced to its corresponding 2D coordinates, maintaining spatial relationships within the scene.
A circular grid centred at the location of the depth sensor is created by rasterising these 2D points. Rasterisation is faster and more efficient than raycasting, making it ideal for real-time processes and handling complex scenes. We do this by calculating the radial distance and azimuthal angle for every point in relation to the sensor. The azimuth angle is derived by calculating the cross-product between the sensor’s directional vector and the vector extending from the sensor to the point of interest.
@@ IV-C Sonification
We demonstrate the ability of the system’s mapping framework to provide contextual information for cognitive mapping and acquiring spatial information. An auditory sensory augmentation framework is applied in which the information from the spatial representation is rendered or sonified as binaural auditory signals. A foundational aspect of the sonification process is to leverage the natural auditory perspective provided by human spatial hearing. In other words, humans naturally perceive the location of sounds around them and we use the mapping framework to take advantage of spatial hearing perception.
@@
In this work, augmented or virtual spatial audio is enabled via the use of binaural room impulse responses (BRIRs) which are filters that mathematically specify the transformation of sound from a particular source location in a room to the ears of the participant. There is a separate BRIR filter for each ear and each source location. As well, BRIR filters are unique for each individual in a perceptually relevant way because of morphological differences in ear shape. The data accumulated from the mapping framework provides a representation of the spatial geometry of the room or environment. This information is sufficient to compute BRIRs using room simulation techniques. However, in this work, we use custom BRIRs that were prerecorded in a representative laboratory environment that includes a polypropylene carpet, a motion capture metallic frame, wooden beams with speakers mounted on them and other furniture (see Fig. 2). Significantly, we restrict the source locations to near-field and far-field locations on the audiovisual horizon because source locations on the audiovisual horizon are perceptually more robust in augmented reality applications using generic, i.e., non-individualized, BRIRs.
@@
The BRIR measurements were recorded using the Brüel & Kjær type 4128C Head and Torso Simulator (HATS) with two in-ear microphones together with a Brüel & Kjær type 9640 turntable. We arranged 10 small loudspeakers in a row at a height consistent with the audiovisual horizon of the HATS model. The 10 loudspeakers were positioned directly in front of the HATS model at zero degree of azimuth at distances varying from 0.4 m to 4 m in steps of 0.4 m. Using the turntable, the HATS model was rotated and BRIRs were recorded every $3^{\circ}$ degrees of azimuth from $-177^{\circ}$ to $180^{\circ}$. We also measured a sensor-to-ear transfer function for the Razer Anzu smart glasses and a headphone-to-ear transfer function for the Sennheiser HD 600 open headphones using the HATS model. Appropriate inverse filters were computed for both the smart glasses and headphones for playback of simulated binaural spatial audio signals. The BRIR files were saved as SOFA (Spatially Oriented Format for Acoustics) files [23].
In order to demonstrate the suitability and ease-of-use of the mapping framework for spatial cognitive mapping, we focus on the 2D circular representation. We demonstrate two sonification modes of operation and various filtering operations. One sonification mode, referred to as circular ranging, sonifies the distance to objects for various directions around the circle. It is ‘as if’ the user reaches out horizontally with an imaginary and magical length-changing cane and swings it in a circle counter clockwise. The cane automatically taps the closest object in each direction as it swings about. Interestingly, the range data from the 2D circular representation is discontinuous. There are jumps at the boundary between different objects because objects are separated and their distances vary randomly. We can use this observation advantageously to create an object-based sonification mode. In other words, we can segment the 2D circular range data at the discontinuities forming ‘pseudo objects’ with an angular width defined by the discontinuities. Each object is then sonified only once in the direction and distance of its nearest point as the imaginary cane swings in a circle. We refer to this sonification mode as circular ranging of objects.
For both sonification modes, various filtering operations can be easily applied, providing ‘sensorisensor’ feedback. In other words, the user controls a fundamental aspect of the sensor system – the filtering operation – and receives sensory feedback in the form of changes in the binaural auditory signal. We now describe several filtering operations:
  • field-of-view sonification filter: during the circular operation, sounds are only generated for specified ranges of azimuthal angle.
  • distance sonification filter: during the circular operation, sounds are only generated for specified ranges of distance.
  • object counting filter: during the “circular ranging of objects”, sounds are only generated for a limited number of objects, controlling object clutter.
@@
The details for the demonstration of the two sonification modes are now described. For circular ranging, the 360-degree of range data are separated into 36 angular sectors spanning $10^{\circ}$. For each angular sector, the angle and range corresponding to the smallest range value are selected. The selected range data for each sector is then converted to binaural spatial audio by applying the appropriate BRIR filter to a ‘tap’ sound chosen from the website freesound.org. Ideally, the ‘tap’ sound would relate to material properties, but we leave this for future work and use a constant ‘tap’ sound in this work. Note that if a particular angular sector has not yet been scanned, a ‘woosh’ sound is played instead of a ‘tap’ sound. The distances encompassing the range data are scaled into the 0.4 m to 4 m range of the BRIRs using a constant scaling chosen appropriately for the given environment. This sonification distance range fits well with the perceptual properties of human distance perception. This is to say, human auditory distance perception is significantly poorer than human auditory direction perception [24] and we generally perceive distance in the near-field ($<$ 2 meters) better than the far-field ($>$ 2 meters). To further enhance the distance perception we apply a slight pitch shift to the tap sound. ‘Tap’ sounds closer than 1.5 m are shifted down in pitch by four semitones and ‘tap’ sounds further than 2.5 m are shifted up in pitch by four semitones. The selection of these parameters are arbitrary and simply chosen for demonstration purposes. For the circular ranging of objects mode, we define a range discontinuity as a change of range, $\Delta$, with $\Delta>0.5$ meters/degree. We also limit the number of objects in a single circular scan to the 5 nearest objects. Again, these parameters are arbitrary and chosen for demonstration purposes.

### v2

# Listen to Your Map: An Online Representation for Spatial Sonification

@@ Abstract
Robotic perception is becoming a key technology for navigation aids, especially helping individuals with visual impairments through spatial sonification. This paper introduces a mapping representation that accurately captures scene geometry for sonification, turning physical spaces into auditory experiences. Using depth sensors, we encode an incrementally built 3D scene into a compact 360-degree representation with angular and distance information, aligning this way with human auditory spatial perception. The proposed framework performs localisation and mapping via VDB-Gaussian Process Distance Fields for efficient online scene reconstruction. The key aspect is a sensor-centric structure that maintains either a 2D-circular or 3D-cylindrical raster-based projection. This spatial representation is then converted into binaural auditory signals using simple pre-recorded responses from a representative room. Quantitative and qualitative evaluations show improvements in accuracy, coverage, timing and suitability for sonification compared to other approaches, with effective handling of dynamic objects as well. An accompanying video demonstrates spatial sonification in room-like environments. [^1] https://tinyurl.com/ListenToYourMap
Keywords: Representation, Sonification, Spatial Perception.
Figure 1: Our sensor-centric representation for spatial sonification. a) 2D circular representation and b) 3D cylindrical representation for the Cow and Lady dataset. The circle is coloured by the distance to the obstacle along each azimuthal angle. We also show the selected close points on the surface. Similarly, for the cylinder, we illustrate the distance up to a certain height.
@@ I Introduction
Robotic perception has the potential to interpret the environment on behalf of humans. Over the past two decades, robotic sensors such as cameras or LiDARs have acted as the eyes of robots, becoming essential tools for gathering information on unknown environments. For individuals living with visual impairments, auditory perception is crucial. This need could be fulfilled by transforming the information accumulated by the robotic perception algorithms via sonification. One of the key steps in representing an environment for sonification is to convert spatial data into an organised format that can be understood through sound. This process facilitates the transformation of physical spaces into auditory experiences, allowing users to understand and navigate through sound. The primary goal of this work is to build a robust framework that can precisely represent an environment’s spatial properties such that it can serve as a basis for sonification.
@@
The ability to take a light-weight sensor information and incrementally build and maintain a compact representation will enable real-time auditory feedback to support spatial awareness and environment interaction, especially for visually impaired individuals. As we want our representation to cater towards sonification for humans, we take motivation from how human beings perceive an environment through sound. Thus, we transform an incrementally built 3D scene captured with depth data (from an RGB-D camera for instance) into a 360-degree projection of distance to the surface information. This approach naturally aligns with human auditory perception while simplifying information to focus on the most relevant spatial points. To achieve this, we propose a mapping representation, based on a sensor-centric organised structure that maintains 2D circular and/or 3D cylindrical rasterised representations (see Fig. 1). These structures capture the geometry of the 3D environment and offer flexibility in how spatial locations are sonified by leveraging VDB-GPDF [1], an online efficient mapping framework based on Gaussian Process (GP) distance and gradient fields and a fast-access VDB data structure (Volumetric, Dynamic grid that shares several characteristics with B+trees [2, 3]). In addition, a localisation framework based on VINS-RGBD [4] is combined with our mapping approach to provide a full online Simultaneous localisation and mapping (SLAM) solution with synchronised poses and raw dense depth data. Moreover, by incorporating custom binaural room impulse responses (BRIRs), our online incrementally-built scene representation offers perceptually robust auditory augmentation of the mapped environment for both near-field and far-field obstacles and unknown space.
@@
We evaluate quantitatively the performance of the proposed mapping representation in terms of accuracy, coverage and timing, demonstrating the ability to convey the required spatial information better than naively using the current sensor information or a state-of-the-art Euclidean Distance Field. In addition, we show qualitative performance in the presence of dynamic objects and benchmarked with the information provided by distance field mapping. Finally, the submitted video showcases the spatial sonification for room-like environments.
@@ II Literature Review
Several studies have explored sonification from vision-based representations. While earlier works focused mostly on 2D vision-to-audio transformations [5], more recent approaches utilise 3D vision-based sensors [6, 7]. Although 3D representations generally offer improved accuracy for localisation tasks, 2D systems can achieve comparable performance for navigation with sonification [8]. Some works have directly used 3D raw data for environment representation [9, 10]. In [9], the authors presented a real-time auditory feedback system that assists blind users in obstacle navigation using a depth camera. The system processes depth data to detect obstacles and sonifies this information. However, using depth data directly can be complex for turning 3D information into clear auditory cues as the information over time can be inconsistent. To tackle this, we propose to reconstruct the environment with sequential depth data and encode it into a consistent 360-degree circular/cylindrical raster-based projection.
@@
More recently, there has been a growing trend towards leveraging machine learning techniques, such as generative models, to enhance environment representation for spatial sonification. These methods allow for more adaptive and context-aware sonification systems, improving the overall user experience. In [11], the authors introduced “SEE-2-SOUND,” a zero-shot framework. They generated spatial audio from visual inputs such as images, GIFs, and videos without pre-training. A similar task was performed in [12], with the additional use of a large language model to generate context-aware synchronised audio-visual content based on the event. While these systems offer adaptability, they often require substantial computational resources. In contrast, our method provides a lightweight, direct mapping representation of spatial data to auditory cues. This ensures real-time responsiveness while preserving essential spatial details and the ability to be deployed on modest embedded devices.
@@
There are different techniques to sonify a virtual environment, Gao et al. showed quantitative evidence of how different sonification strategies can influence user perception and behaviour in [13]. They compared four elevation sonification mapping strategies and found that binary relative mapping performs better than the others, and in combination with azimuth sonification, it gives clear directional information generating more accurate and efficient task completion. Keeping this in mind, our proposed representation technique is designed to effectively capture both angular and distance information. Thus, ensuring optimal translation of spatial data into auditory cues for the sonification process.
@@
There have been many works on sonification that are dedicated towards visually impaired persons[14]. The type of sound used for sonification also varies depending on the task and the type of information it carries. While discrete sounds are good for simpler information and objects that are near the user, continuous sounds are more suited to deliver complex elements in the scene [15]. Schwartz et al. proposed a mobile application that can reconstruct and sonify a scene to guide users with visual impairment in real-time [14]. The work in [16] built a 3D sonification framework to sonify images from video sequences to send the user a predefined sonic sensorial description through bone-conduction headphones. They combined YOLOV3 [17] for object detection with a probabilistic occupancy map from OctoMap [18] to detect environmental anomalies. These works primarily focus on sonifying immediate surroundings or predefined objects with limited adaptability. They do not offer a full 360-degree spatial coverage, focusing mostly on near-field obstacles. Our proposed approach, in contrast, is flexible, robust and offers greater adaptability as it conveys only spatial information for near-field and far-field.
@@ III Preliminaries
III-A Localisation
We employ a modified VINS-RGBD [4] localisation framework to generate precise camera poses as inputs for the mapping. A key aspect of our approach is the improved synchronisation between the estimated poses and the raw dense point cloud data captured by the RGB-D camera. By refining the VINS framework, we ensure consistent spatial alignment, which enhances the accuracy and reliability of the mapping process, ultimately enabling more precise and detailed scene reconstruction.
III-B VDB-GPDF Mapping
@@
For testing, a set of query points is generated by ray-casting from the sensor origin through the voxels in the current reference frame. Each test point is used to query the L-GPDF for the distance field and variance estimates. The queried values from L-GPDF are then fused into a global VDB grid map by updating the voxel distances using the weighted sum method. Following the fusion of the distances and surface properties, the marching cubes algorithm [22] is used to reconstruct a dense surface by identifying the active voxels in the global VDB. The Global Gaussian Process Distance Field (G-GPDF) is made up of the zero-crossing points from active leaf nodes. At this stage, queries for the G-GPDF are computed by averaging the inferred distances from neighbouring GP nodes. For further details refer to [1].
IV Circular/Cylindrical Rasterisation
We aim to incrementally build a scene representation that facilitates spatial sonification of an unknown environment. We propose a representation based on VDB-GPDF that is encoded to cater for the sonification requirements of future human sensory augmentation. Given the sequentially built G-GPDF model and the extracted 3D surface, we encode the scene into a 2D or 3D 360-degree raster-based projection. With this representation, pre-recorded binaural room impulse response filters are used to create the rendered spatial sound.
@@ IV Circular/Cylindrical Rasterisation
Spatial audio feedback requires accurate but simple information, thus we propose to represent the incrementally built dense 3D map of the environment as a 2D circular sensor-centric representation. To ensure our representation aligns with human navigation requirements, we segment the ground and non-ground points. Any point above the ground and in the height range of a human that can potentially be an obstacle is classified as a non-ground point. Then, we project all these 3D non-ground points onto a 2D plane. On this plane, every 3D point is reduced to its corresponding 2D coordinates, maintaining spatial relationships within the scene.
A circular grid centred at the location of the depth sensor is created by rasterising these 2D points. Rasterisation is faster and more efficient than raycasting, making it ideal for real-time processes and handling complex scenes. We do this by calculating the radial distance and azimuthal angle for every point with respect to the sensor. The azimuthal angle is derived by calculating the cross-product between the sensor’s directional vector and the vector extending from the sensor to the point of interest.
@@ IV-C Sonification
We demonstrate the ability of our proposed representation to provide spatial information for cognitive sonification. An auditory framework is applied with the spatial representation to generate binaural auditory signals. A foundational aspect of the sonification process is to leverage the natural auditory perspective provided by human spatial hearing. In other words, humans naturally perceive the location of sounds around them and we design the representation to take advantage of spatial hearing perception.
@@
In this work, augmented or virtual spatial audio is enabled via the use of binaural room impulse responses (BRIRs) which are filters that mathematically specify the transformation of sound from a particular source location in a room to the ears of the individual listening to the map. There is a separate BRIR filter for each ear and each source location. Note that BRIR filters are unique for each individual in a perceptually relevant way because of morphological differences in ear shape. The data accumulated from the map representation has the spatial geometry of the room or environment. This information is sufficient to compute BRIRs using room simulation techniques. However, in this work, we use custom BRIRs that were prerecorded in a representative laboratory environment that includes a polypropylene carpet, a motion capture metallic frame, wooden beams with speakers mounted on them and other furniture (see Fig. 2).
@@
The BRIR measurements were recorded using the Brüel & Kjær type 4128C Head and Torso Simulator (HATS) with two in-ear microphones together with a Brüel & Kjær type 9640 turntable. We arranged 10 small loudspeakers in a row at a height consistent with the audiovisual horizon of the HATS model. The 10 loudspeakers were positioned directly in front of the HATS model at zero degree of azimuth at distances varying from 0.4 m to 4 m in steps of 0.4 m. Using the turntable, the HATS model was rotated and BRIRs were recorded every $3^{\circ}$ of azimuth from $-177^{\circ}$ to $180^{\circ}$. The BRIR files were saved as Spatially Oriented Format for Acoustics (SOFA) files [23]. The full loudspeaker recording positions are shown in Fig. 3.
In order to demonstrate the suitability of the mapping representation for spatial sonification, we focus on the 2D circular representation. We sonify the distance to the surface for all directions around the circle counter-clockwise. It simulates a scanning “length-changing cane”. The cane automatically taps the closest surface in each discrete direction as it scans. As we mentioned, the BRIRs were recorded every $3^{\circ}$. Note that, our circular and cylindrical representations have $1^{\circ}$ resolution. To avoid sensory overload, we choose $10^{\circ}$ as the resolution to separate the 360-degree range data into 36 angular sectors.
@@
For each angular sector, the angle and range corresponding to the smallest range value are selected. The selected range data for each sector is then converted to binaural spatial audio by applying the appropriate BRIR filter to a ‘tap’ sound. Note that if a particular angular sector has not yet been scanned, a ‘woosh’ sound is played instead of a ‘tap’ sound. The distances encompassing the range data are scaled into the 0.4 m to 4 m range of the BRIRs using a constant scaling chosen appropriately for the given environment. Distances beyond 4 m are clamped to 4 m in the BRIRs range. This sonification distance range fits well with the perceptual properties of human distance perception, which is significantly poorer than human auditory direction perception [24] and we generally perceive distance in the near-field ($<$ 2 meters) better than the far-field ($>$ 2 meters). To further enhance the distance perception we apply a slight pitch shift to the tap sound. ‘Tap’ sounds closer than 1.5 m are shifted down in pitch by four semitones and ‘tap’ sounds further than 2.5 m are shifted up in pitch by four semitones. Our accompanying video shows these effects. Listening through headphones is recommended to hear the sounds at the correct azimuth location.